# Gravitational Wave Memory from Accelerating Relativistic Jets in Multiple Thick Shell Scenarios

### v1

@@ Abstract
Gravitational wave (GW) memory, a permanent distortion of the space-time metric, is anticipated during the acceleration of relativistic jets in gamma-ray bursts (GRBs). While the precise mechanism behind GRBs is not yet fully understood, detecting GW memory may contribute to clarifying their nature. In this paper, we consider various scenarios of GW memory emission, including both single and multiple shells with thin- and thick-shells. In particular, the memory spectrum for each scenario is compared with the sensitivity of next-generation detectors, namely DECIGO and ET-D. Physical properties spread over a broad-band region, emphasizing the importance of combined and wide-band observations. We also simulate GW memory based on nearby, realistic scenarios and demonstrate its detectability.
@@ I Introduction
For a GRB jet with the bulk Lorentz factor greater than $\sim 10^{2}$, the jet acceleration involving a large mass may emit detectable GW memories [2, 3, 4, 5, 6]. In such cases, the metric perturbation due to the GW permanently deviates from its original value and remains. This imprint is known as the GW memory effect. The GW memory provides information about the jet structure, acceleration processes, and the history of the mass ejection from the central engine. Relativistic jets may also arise in core-collapse supernovae [7], binary neutron star mergers [8], tidal disruption events [9], and even in giant flares of magnetars [10]. Some of these jets are choked during propagation through the dense surrounding medium, and in such a case, a bright electromagnetic signal is unlikely. Therefore, the detection of such GW memory induced by these astrophysical phenomena will give us not only the evidence for the relativistic jets but also the mechanism of such high-energy astrophysical phenomena. Note that in this paper, we only consider baryonic matter-dominated jets accelerating to ultra-relativistic speeds, while GW memory effect may also arise from non-baryonic components, such as neutrino-driven jets [11] and also non-relativistic aspherical supernovae [12, 13, 14].
@@
Theoretical studies on GW memory emission during the initial jet acceleration phase have progressed as follows. Segalis and Ori [2] considered GW memory induced by a point particle that was ejected from the central engine and accelerated to ultra-relativistic speed. Sago et al. [4] extended a formalism for an infinitesimally thin shell with finite angular size, expressing the collimated jet. They simulated emission from multiple shells using step functions as the expression of GW memory, meaning each shell accelerates instantaneously. Birnholtz and Piran [5] relaxed the assumption of instantaneous acceleration, and calculated the signal from the continuously accelerating axisymmetric jet. Furthermore, Leiderschneider and Piran [6] considered more generalized cases in which prolonged acceleration and prolonged duration of ejection of the jet from the central engine.
@@
After the initial acceleration, the jet propagates into the dense surroundings such as progenitor star in long GRBs or merger ejecta in short GRBs, and changes its velocity, resulting in an additional shift of the metric perturbations. Using thin shell approximation, Yu [15] calculated GW memory from the jet interacting with the star’s envelope or the ejecta. Moreover, Urrutia et al. [16] performed hydrodynamic simulations of jets that penetrate collapsing stars. In their simulation, the pre-accelerated jet is injected at the inner boundary at $5\times 10^{8}$ cm from the source, with an initial Lorentz factor of 10. The time evolution of the GW memory is derived, highlighting a double peak in the time domain; one originates from the initial jet acceleration to the final Lorentz factor, and the other is from the jet-stellar envelope interaction near the low-density surface envelope of the star. Later, the jet breaks out the dense material and begins to emit gamma-rays. The dissipation of the jet’s kinetic energy into photons occurs far from the central engine ($\gtrsim 10^{13}$ cm), at which again causes the variation of the metric perturbation. Taking into account the photon emission, Akiba et al. [17] calculated the GW memory from the multiple massive points colliding with each other. They were interested in the simultaneous detection of gamma-ray photons and GW memory so that the jet was viewed on-axis; that is, the angle between the jet axis and the observer’s line of sight was less than the inverse of the bulk Lorentz factor of the jet. In such a case, the anti-beaming effect [5, 4] reduces the observed GW amplitude. Note that detecting the GW memory requires the source to be nearby, which increases the likelihood of viewing the jet off-axis. However, the GW memory may still be detectable due to the anti-beaming effect. For nearby off-axis viewing events, observed photon flux is still detectable despite of the relativistic beaming effect [18, 19, 20, 21, 22, 23]. Moreover, the GW amplitude can be larger for the off-axis cases than the on-axis cases. Finally, in Huang et al. [24] the GW memory from a thin-shell jet that receives late-time energy injection to account for the observed shallow decay phase in the canonical X-ray afterglow [25].
@@
Gamma-ray light curves for long GRB consist of many pulses (see Ref. [26] for a review). According to the internal shock model of the prompt gamma-ray emission [27, 28], the density in the jet is highly inhomogeneous in the radial direction, and the high-density regions, which we refer to as shells hereafter, have different velocities, so that the shells collide with each other, causing shocks that accelerate gamma-ray emitting high-energy electrons. A collision of two shells produces a pulse in gamma-ray light curves. Since typical long GRBs have around $10^{2}$ pulses, the jets are thought to contain approximately $10^{2}$ shells. The finite rising time scale of the gamma-ray pulse reflects the lifetime of the internal shocks propagating into the shell with finite thickness. Taking into account the internal shock model, Akiba et al. [17] calculated GW memory in the point mass approximation. In this paper, extending previous theoretical works [4, 5, 6, 17], we consider the GW memory from an accelerating jet that consists of many thick shells. In Sago et al. [4], the jet consists of many patchy, infinitesimally thin shells accelerating instantaneously. Birnholtz and Piran [5] considered GW memory emission from an angularly structured thin shell. In Leiderschneider and Piran [6], various cases are considered, one of which is a jet with a thick shell formed by continuous mass ejection. The present work builds upon the study conducted in [6], with a particular emphasis on the role of radial inhomogeneity. We calculate the observed spectrum of GW memory in specific cases and discuss its detectability by next-generation detectors, such as DECIGO [29] and Einstein Telescope (ET-D) [30].
@@ II Waveform formulation
According to the internal shock model of the prompt GRB emission, we consider an accelerating jet with $N$ thick shells. We take the spherical coordinate system $(t,r,\theta,\phi)$ in the rest frame of the central engine. As shown in Fig. 1, each shell shares a common opening half-angle $\Delta\theta$ and the viewing angle between the observer and the jet central axis $\theta_{v}$, where $\theta=0$ means the direction toward the observer. The shells emerge at $r=r_{0}$ and subsequently begin to accelerate. In the following, we measure the Lorentz factor of the moving shells, $\gamma=1/\sqrt{1-\beta^{2}}$, where $\beta$ is their velocity divided by the speed of light $c$, in the central engine frame.
We approximate the GW memory emission of a thick shell as the superposition of emissions $n$ infinitesimally thin shells, with sufficiently large $n$ (finally, we apply $n\to\infty$), Hence, before addressing the case of thick shells, we will elaborate on the multiple thin shell case.
@@ II.1 Single / Multiple thin shells
in the case of ultra-relativistic limit ($\gamma_{j}\gg 1$) and small $\theta$.
@@
where $\beta(\gamma)=\sqrt{1-1/\gamma^{2}}$ hereafter, $\Delta\Omega=2\pi(1-\cos\Delta\theta)$ is the solid angle of the shell. In this context, a point mass is calculated as $m_{j}/\Delta\Omega$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
where the energy of $j$-th shell is defined as:
@@
The Fourier component of the GW memory from multiple thin shells is calculated by the superposition of the GW memory from $N$ shells, incorporating time-shifts of the Fourier component for each shell as follows:
@@ III Waveform properties
In this section, we compute GW memories under various conditions. We investigate the differences in the representation of the GW memory waveform when the thickness $W_{j}$, the terminal Lorentz factor $\gamma_{j}$, the number of shells $N$, the time interval between shells $\delta\tau_{j}=\tau_{j+1}-\tau_{j}$ ($j=1,\cdots,N-1$) vary. The total time in a central engine emitting GW memories is denoted as:
@@
We use the following configuration: the opening half-angle and the viewing angle are set to $\Delta\theta=\theta_{v}=0.1$ rad, and the starting position for acceleration is $r_{0}=10^{7}~{}\mathrm{cm}$, except for Sec. III.1. The distance to the source is also set to $R=10~{}\mathrm{Mpc}$ in this section.
@@ III.1.1 Dependency of $\Delta\theta$ and $\theta_{v}$
We consider GW memory from a single thin shell using Eq. (11) under conditions $N=1$ and $\gamma_{1}=100$. Figure 3 shows the Fourier waveform with a different opening half-angle $\Delta\theta$ and a viewing angle $\theta_{v}$. Their total energies are the same as $\mathcal{E}=10^{52}~{}\mathrm{erg}$ for comparison. The waveform with $\theta_{v}=0.1$ rad as shown in Fig. 3 indicates that their amplitude depends on the opening half-angle $\Delta\theta$. In the case where $\Delta\theta\leq\theta_{v}$, represented by the dotted blue and solid purple lines, the situation is considered where the jet’s central axis is far from the observer, and the amplitude remains almost unchanged. On the other hand, for $\Delta\theta>\theta_{v}$, represented by the dashed orange and dot-dashed green lines, the jet’s central axis gradually approaches the observer, and their amplitude decreases.
Figure 3 shows the opposite situation, where $\theta_{v}$ is varied while keeping $\Delta\theta$ constant. The dashed red line, $\theta_{v}=0.02$, represents the case where the observer is closest to the jet’s central axis, and its amplitude is the smallest. This phenomenon, in which the amplitude decreases across the entire frequency range, is known as the anti-beaming effect [3] and was demonstrated by Sago et al. [4] in the context of jetted GW memory emission. Although our waveform formula given in Eq. (11) includes an additional dependence on the $T_{\mathrm{obs}}(\theta,\gamma_{j})$ term, we still confirm the presence of the anti-beaming effect. Furthermore, our waveform for a single thin shell is consistent with previous studies[4, 5, 6]. As $\theta_{v}$ increases, the initial attenuation appears in the low-frequency range, such as around 100 Hz when $\theta_{v}=0.7$. This is due to the contribution of $\cos\theta$ in Eq. (2) which becomes non-negligible. Hereafter, we set $\Delta\theta=\theta_{v}=0.1$ rad in this paper to simplify the discussion.
@@ III.1.2 Dependency of the initial radius $r_{0}$
Figure 4: The waveforms for different values of the initial radius $r_{0}$. The solid black line, which is the same as the solid purple lines ($\Delta\theta=\theta_{v}=0.1~{}\mathrm{rad}$) in Fig. 3, is plotted as a reference. The total energy, the distance to the source, final bulk Lorentz factor are chosen as $\mathcal{E}=10^{52}~{}\mathrm{erg}$, $R=10~{}\mathrm{Mpc}$, and $\gamma_{1}$ = 100, respectively.
An order of magnitude for the observed acceleration time $T_{\mathrm{obs}}$, given by Eq. (3), is mainly determined by the initial radius, $r_{0}$. We compute waveforms for several values of $r_{0}$ as shown in Fig. 4. The low-frequency components at $f\ll 1$ Hz behave as the same amplitude. At low frequencies, where the second and higher-order terms in the Taylor expansion of the exponential function are ignored, the characteristic strain in Eq. (8) is expressed as:

### v2

@@ Abstract
Gravitational wave (GW) memory, a permanent distortion of the space-time metric, is anticipated during the acceleration of relativistic jets in gamma-ray bursts (GRBs). While the precise mechanism behind GRBs is not yet fully understood, detecting GW memory may contribute to clarifying their nature. In this paper, we consider various scenarios of GW memory emission, including both single and multiple shells with thin and thickshells. In particular, the memory spectrum for each scenario is compared with the sensitivity of next-generation detectors, namely Deci-hertz Interferometer Gravitational-Wave Observatory and Einstein Telescope. Physical properties spread over a broad-band region, emphasizing the importance of combined and wide-band observations. We also simulate GW memory based on nearby, realistic scenarios and demonstrate its detectability.
@@ I Introduction
For a GRB jet with the bulk Lorentz factor greater than $\sim 10^{2}$, the jet acceleration involving a large mass may emit detectable GW memories [2, 3, 4, 5, 6, 7]. In such cases, the metric perturbation due to the GW permanently deviates from its original value and remains. This imprint is known as the GW memory effect. The GW memory provides information about the jet structure, acceleration processes, and the history of the mass ejection from the central engine. Relativistic jets may also arise in core-collapse supernovae [8], binary neutron star mergers [9], tidal disruption events [10], and even in giant flares of magnetars [11, 7]. Some of these jets are choked during propagation through the dense surrounding medium, and in such a case, a bright electromagnetic signal is unlikely. Therefore, the detection of such GW memory induced by these astrophysical phenomena will give us not only the evidence for the relativistic jets but also the mechanism of such high-energy astrophysical phenomena. Note that in this paper, we only consider baryonic matter-dominated jets accelerating to ultrarelativistic speeds, while the GW memory effect may also arise from nonbaryonic components, such as neutrino-driven jets [12] and also nonrelativistic aspherical supernovae [13, 14, 15].
@@
Theoretical studies on GW memory emission during the initial jet acceleration phase have progressed as follows. Segalis and Ori [2] considered GW memory induced by a point particle that was ejected from the central engine and accelerated to ultrarelativistic speed. Sago et al. [4] extended a formalism for an infinitesimally thin shell with finite angular size, expressing the collimated jet. They simulated emission from multiple shells using step functions as the expression of GW memory, meaning each shell accelerates instantaneously. Birnholtz and Piran [5] relaxed the assumption of instantaneous acceleration, and calculated the signal from the continuously accelerating axisymmetric jet. Furthermore, Leiderschneider and Piran [6] considered more general cases involving prolonged acceleration and prolonged jet ejection from the central engine.
@@
After the initial acceleration, the jet propagates into the dense surroundings such as the progenitor star in long GRBs or merger ejecta in short GRBs, and changes its velocity, resulting in an additional shift of the metric perturbations. Using thin shell approximation, Yu [16] calculated GW memory from the jet interacting with the star’s envelope or the ejecta. Moreover, Urrutia et al. [17] performed hydrodynamic simulations of jets that penetrate collapsing stars. In their simulation, the preaccelerated jet is injected at the inner boundary at $5\times 10^{8}$ cm from the source, with an initial Lorentz factor of 10. The time evolution of the GW memory is derived, highlighting a double peak in the time domain; one originates from the initial jet acceleration to the final Lorentz factor, and the other is from the jet-stellar envelope interaction near the low-density surface envelope of the star. Later, the jet breaks out the dense material and begins to emit gamma rays. The dissipation of the jet’s kinetic energy into photons occurs far from the central engine ($\gtrsim 10^{13}$ cm), which again causes the variation of the metric perturbation. Taking into account the photon emission, Akiba et al. [18] calculated the GW memory from the multiple massive points colliding with each other. They were interested in the simultaneous detection of gamma-ray photons and GW memory so that the jet was viewed on axis; that is, the angle between the jet axis and the observer’s line of sight was less than the inverse of the bulk Lorentz factor of the jet. In such a case, the antibeaming effect [5, 4] reduces the observed GW amplitude. Note that detecting the GW memory requires the source to be nearby, which increases the likelihood of viewing the jet off axis. However, the GW memory may still be detectable due to the antibeaming effect. For nearby off axis viewing events, observed photon flux is still detectable despite of the relativistic beaming effect [19, 20, 21, 22, 23, 24]. Moreover, the GW amplitude can be larger for the off axis cases than the on axis cases. Finally, in Huang et al. [25] studied the GW memory from a thin-shell jet with late-time energy injection to account for the observed shallow decay phase in the canonical X-ray afterglow [26].
@@
Gamma-ray light curves for long GRB consist of many pulses (see Ref. [27] for a review). According to the internal shock model of the prompt gamma-ray emission [28, 29], the density in the jet is highly inhomogeneous in the radial direction, and the high-density regions, which we refer to as shells hereafter, have different velocities, so that the shells collide with each other, causing shocks that accelerate gamma-ray emitting high-energy electrons. A collision of two shells produces a pulse in gamma-ray light curves. Since typical long GRBs have around $10^{2}$ pulses, the jets are thought to contain approximately $10^{2}$ shells. The finite rising timescale of the gamma-ray pulse reflects the lifetime of the internal shocks propagating into the shell with finite thickness. Taking into account the internal shock model, Akiba et al. [18] calculated GW memory in the point mass approximation. In this paper, extending previous theoretical works [4, 5, 6, 18], we consider the GW memory from an accelerating jet that consists of many thick shells. In Sago et al. [4], the jet consists of many patchy, infinitesimally thin shells accelerating instantaneously. Birnholtz and Piran [5] considered GW memory emission from an angularly structured thin shell. In Leiderschneider and Piran [6], various cases are considered, one of which is a jet with a thick shell formed by continuous mass ejection. The present work builds upon the study conducted in [6], with a particular emphasis on the role of radial inhomogeneity. We calculate the observed spectrum of GW memory in specific cases and discuss its detectability by next-generation detectors, such as Deci-hertz Interferometer Gravitational-Wave Observatory (DECIGO) [30] and the Einstein Telescope (ET-D) [31].
@@ II Waveform formulation
According to the internal shock model of the prompt GRB emission, we consider an accelerating jet with $N$ thick shells. We take the spherical coordinate system $(t,r,\theta,\phi)$ in the rest frame of the central engine. As shown in Fig. 1, each shell shares a common opening half angle $\Delta\theta$ and the viewing angle between the observer and the jet central axis $\theta_{v}$, where $\theta=0$ means the direction toward the observer. The shells emerge at $r=r_{0}$ and subsequently begin to accelerate. In the following, we measure the Lorentz factor of the moving shells, $\gamma=1/\sqrt{1-\beta^{2}}$, where $\beta$ is their velocity divided by the speed of light $c$, in the central engine frame.
We approximate the GW memory emission of a thick shell as the superposition of emission $n$ infinitesimally thin shells, with sufficiently large $n$ (finally, we apply $n\to\infty$), Hence, before addressing the case of thick shells, we will elaborate on the multiple thin shell case.
@@ II.1 Single / multiple thin shells
in the case of the ultrarelativistic limit ($\gamma_{j}\gg 1$) and small $\theta$.
@@
where $\beta(\gamma)=\sqrt{1-1/\gamma^{2}}$ hereafter, and $\Delta\Omega=2\pi(1-\cos\Delta\theta)$ is the solid angle of the shell. In this context, a point mass is calculated as $m_{j}/\Delta\Omega$.
@@
$\eta(T,\theta,\gamma_{j})=\begin{cases}\ 0&(T<T_{s}(\theta)),\\ \ (T-T_{s}(\theta))/{T_{\mathrm{obs}}(\theta,\gamma_{j})}&(T_{s}(\theta)<T<T_{\mathrm{E}}(\theta,\gamma_{j})),\\ \ 1&(T>T_{\mathrm{E}}(\theta,\gamma_{j}))~~.\end{cases}$ (5)
@@
$h_{\mathrm{thin}}(T,\gamma_{j})=\int h_{\mathrm{point}}(T,\theta,\phi,\gamma_{j})d\Omega=\frac{2\gamma_{j}\beta^{2}(\gamma_{j})m_{j}G}{c^{2}R\Delta\Omega}\int^{\Delta\theta+\theta_{v}}_{\max{[0,\Delta\theta-\theta_{v}]}}\frac{\sin^{3}\theta\sin(2\Delta\phi)}{1-\beta\cos\theta}\eta(T,\theta,\gamma_{j})\ d\theta,$ (6)
@@
$\Delta\phi=\begin{cases}\pi:\Delta\theta>\theta_{v}\>\>\mathrm{and}\>\>0<\theta\leq\Delta\theta-\theta_{v},\\ \arccos\left(\cfrac{\cos\Delta\theta-\cos\theta_{v}\cos\theta}{\sin\theta_{v}\sin\theta}\right):\mathrm{otherwise}.\end{cases}$ (7)
@@
$\begin{split}\tilde{h}_{\mathrm{thin}}(f,\gamma_{j})&=\frac{2\beta^{2}(\gamma_{j})\mathcal{E}_{j}G}{c^{4}R\Delta\Omega}\int^{\theta_{v}+\Delta\theta}_{\max[0,\theta_{v}-\Delta\theta]}\frac{\sin^{3}\theta\sin(2\Delta\phi)}{1-\beta(\gamma_{j})\cos\theta}\frac{(e^{-i2\pi fT_{\mathrm{obs}}(\theta,\gamma_{j})}-1)}{(2\pi f)^{2}T_{\text{obs}}(\theta,\gamma_{j})}e^{-i2\pi fT_{s}(\theta)}\ d\theta,\end{split}$ (8)
where the energy of the $j$-th shell is defined as
@@
The Fourier component of the GW memory from multiple thin shells is calculated by the superposition of the GW memory from $N$ shells, incorporating time shifts of the Fourier component for each shell as follows:
@@ II.2 Single / multiple thick shells
In this paper, we adopt a shell thickness $W_{j}$ in the range, from $10^{7}$ to $10^{9}$ cm [1, 27]. The typical value of the observationally estimated collimation-corrected gamma-ray energy, that is the gamma-ray emission energy confined in the jet, is about $10^{50-51}$ erg [36]. Assuming a gamma-ray emission efficiency, that is, the conversion efficiency from jet kinetic energy to gamma-ray energy, of about 5–10%, we estimate the jet kinetic energy to be on the order, $\mathcal{E}\sim 10^{52}$ erg, that is the value adopted in the following of this paper.
@@ III Waveform properties
In this section, we compute GW memories under various conditions. We investigate the differences in the representation of the GW memory waveform when the thickness $W_{j}$, the terminal Lorentz factor $\gamma_{j}$, the number of shells $N$, and the time interval between shells $\delta\tau_{j}=\tau_{j+1}-\tau_{j}$ ($j=1,\cdots,N-1$) vary. The total time in a central engine emitting GW memories is denoted as:
@@
We use the following configuration: the opening half angle and the viewing angle are set to $\Delta\theta=\theta_{v}=0.1$ rad, and the starting position for acceleration is $r_{0}=10^{7}~\mathrm{cm}$, except for Sec. III.1. The distance to the source is also set to $R=10~\mathrm{Mpc}$ in this section.
@@ III.1.1 Dependency of $\Delta\theta$ and $\theta_{v}$
We consider GW memory from a single thin shell using Eq. (11) under conditions $N=1$ and $\gamma_{1}=100$. Figure 3 shows the Fourier waveform with a different opening half angle $\Delta\theta$ and a viewing angle $\theta_{v}$. Their total energies are the same as $\mathcal{E}=10^{52}~\mathrm{erg}$ for comparison. The waveform with $\theta_{v}=0.1$ rad as shown in Fig. 3 indicates that their amplitude depends on the opening half angle $\Delta\theta$. In the case where $\Delta\theta\leq\theta_{v}$, represented by the dotted blue and solid purple lines, the situation is considered where the jet’s central axis is far from the observer, and the amplitude remains almost unchanged. On the other hand, for $\Delta\theta>\theta_{v}$, represented by the dashed orange and dot-dashed green lines, the jet’s central axis gradually approaches the observer, and their amplitude decreases.
Figure 3 shows the opposite situation, where $\theta_{v}$ is varied while keeping $\Delta\theta$ constant. The dashed red line, $\theta_{v}=0.02$, represents the case where the observer is closest to the jet’s central axis, and its amplitude is the smallest. This phenomenon, in which the amplitude decreases across the entire frequency range, is known as the antibeaming effect [3] and was demonstrated by Sago et al. [4] in the context of jetted GW memory emission. Although our waveform formula given in Eq. (11) includes an additional dependence on the $T_{\mathrm{obs}}(\theta,\gamma_{j})$ term, we still confirm the presence of the antibeaming effect. Furthermore, our waveform for a single thin shell is consistent with previous studies[4, 5, 6]. As $\theta_{v}$ increases, the initial attenuation appears in the low-frequency range, such as around 100 Hz when $\theta_{v}=0.7$. This is due to the contribution of $\cos\theta$ in Eq. (2) which becomes non-negligible. Hereafter, we set $\Delta\theta=\theta_{v}=0.1$ rad in this paper to simplify the discussion.
@@ III.1.2 Dependency of the initial radius $r_{0}$
Figure 4: The waveforms for different values of the initial radius $r_{0}$. The solid black line, which is the same as the solid purple lines ($\Delta\theta=\theta_{v}=0.1~\mathrm{rad}$) in Fig. 3, is plotted as a reference. The total energy, the distance to the source, and the final bulk Lorentz factor are chosen as $\mathcal{E}=10^{52}~\mathrm{erg}$, $R=10~\mathrm{Mpc}$, and $\gamma_{1}$ = 100, respectively.
An order of magnitude for the observed acceleration time $T_{\mathrm{obs}}$, given by Eq. (3), is mainly determined by the initial radius, $r_{0}$. We compute waveforms for several values of $r_{0}$ as shown in Fig. 4. The low-frequency components at $f\ll 1$ Hz behave as the same amplitude. At low frequencies, where the second- and higher-order terms in the Taylor expansion of the exponential function are ignored, the characteristic strain in Eq. (8) is expressed as: